OpenStudy (anonymous):

Can someone please solve 15^2+b^2=17^2 !!!!!!!!!!!!!

OpenStudy (anonymous):

15*15=? +1*1=?

OpenStudy (anonymous):

You get it?

OpenStudy (anonymous):

so the answer would be 8 right?

OpenStudy (anonymous):

Where do you get 8 from ?

OpenStudy (anonymous):

225+b^2=289 b^2=64 sq rt both sides b=8

OpenStudy (anonymous):

Correct (:

OpenStudy (anonymous):

thanks(: